OpenStudy (anonymous):

5-15). Multiply the following binomials using the FOIL or the COLUMN methods. (x+5) (x+7)

OpenStudy (anonymous):

x^2+12x+35

OpenStudy (anonymous):

The method is x(x+5) + 7(x+5) = x^2 + 5x + 7x + 35 = x^2 + 12x + 35